### v1

# Quantization Effects on Neural Networks Perception: How would quantization change the perceptual field of vision models?

Abstract
Neural network quantization is an essential technique for deploying models on resource-constrained devices. However, its impact on model perceptual fields, particularly regarding class activation maps (CAMs), remains a significant area of investigation. In this study, we explore how quantization alters the spatial recognition ability of the perceptual field of vision models, shedding light on the alignment between CAMs and visual saliency maps across various architectures. Leveraging a dataset of 10,000 images from ImageNet, we rigorously evaluate six diverse foundational CNNs: VGG16, ResNet50, EfficientNet, MobileNet, SqueezeNet, and DenseNet. We uncover nuanced changes in CAMs and their alignment with human visual saliency maps through systematic quantization techniques applied to these models. Our findings reveal the varying sensitivities of different architectures to quantization and underscore its implications for real-world applications in terms of model performance and interpretability. The primary contribution of this work revolves around deepening our understanding of neural network quantization, providing insights crucial for deploying efficient and interpretable models in practical settings.
Index Terms—  Neural Network Quantization, Class Activation Maps (CAM), Model Interpretability, Visual Saliency
1 Introduction
The rapid advancements in deep learning have led to the development of increasingly complex neural network architectures. These models have demonstrated remarkable performance across various tasks, such as image classification, object detection, and segmentation. However, deploying these models on resource-constrained devices remains challenging due to their high computational and memory requirements.
Neural network quantization [1] has emerged as a key technique for addressing the deployment challenges by reducing model size and computational complexity. Quantization involves converting the weights and activations of a deep neural network from floating-point precision to lower bit-width representations. Quantization techniques all evolved from the following equation :
$Q(x)=\text{round}\left(\frac{x}{\Delta}\right)\times\Delta\vspace{-2mm}$ (1)
where $x$ is the continuous value, $Q(x)$ presents the quantized value, and $\Delta$ is the quantization step size, determining the precision of the quantized representation.
By reducing the precision of parameters, quantization enables the delivery of lighter-weight models for efficient execution on devices with limited computational resources, such as smartphones [2], embedded systems [3], and IoT devices [4].
While quantization offers practical benefits, it introduces changes to model behavior that may impact interpretability. Interpretable models are crucial for understanding the decision-making process of neural networks, particularly in safety-critical applications such as medical diagnosis and autonomous driving. Ensuring that quantized models maintain interpretability is essential for building trust and facilitating human-machine collaboration.
Class activation maps (CAMs) [5] are visualization techniques that highlight the regions of an image contributing most to a neural network’s decision. CAMs provide valuable insights into where the model focuses its attention when making predictions, enhancing understanding and interoperability. Some algorithms like Grad-CAM [6] and Grad-CAM++[7] analyze the gradients flowing back into the final convolutional layer of a CNN. These gradients highlight which features in the image contributed most to the model’s decision, and the algorithm uses them to create a CAM pinpointing these important regions. Understanding how quantization affects CAMs can provide insights into the changes in model behavior and decision-making processes.
Visual saliency [8] refers to the perceptual prominence or importance of different regions within an image for the human visual system. Unlike CAMs, which are derived from the internal activations of a neural network, visual saliency is based on human perception and attention. Comparing CAMs with visual saliency maps [8] can help assess the alignment between model attention and human perception and evaluate the interpretability of quantized models.
In this paper, we investigate the impact of quantization on model perception, focusing on CAMs and their alignment with visual saliency maps. We evaluate six diverse neural network architectures on a comprehensive dataset of images from ImageNet [9]. Through rigorous experimentation, we analyze the changes in CAMs induced by quantization and their implications for model interpretability and performance.  [^1] Source Code
Our findings contribute to advancing the understanding of neural network quantization and offer insights crucial for deploying efficient and interpretable models in practical settings.
2 Methodology
Fig. 1: The experimental methodology of experimentation.
In this study, we aim to explore the effects of quantization on neural network model perception. By investigating how quantization affects CAMs across different architectures, we hope to uncover insights that can enhance model efficiency and interpretability in real-world applications.
2.1 Model Selection and Quantization
We selected six widely-used neural network architectures, namely VGG16[10], ResNet50[11], EfficientNet[12], MobileNet[13], SqueezeNet[14], and DenseNet[15], due to their diverse structures and widespread adoption across various computer vision tasks.
Quantization, the focal subject of our study, was performed using the Quantization Aware Training (QAT) module from PyTorch 222 Quantization Pytorch only for inference, in order to simulate adaptive quantization. This technique allows for the quantization of neural network weights and activations while preserving the gradient flow necessary for calculating CAMs. Three Quantization levels were chosen: namely Float 32 (f32), Integer 16 (int16), and Integer 8 (int8). It’s worth noting that other quantization techniques in PyTorch do not support gradient flow during quantization.
2.2 CAM Calculation
CAMs were computed for each image in the subset chosen from the validation set of ImageNet [9] across different quantization levels. CAMs provide insights into the regions of an image that contribute most to a neural network’s decision-making process, facilitating interpretability and understanding of model behavior.
The CamGrad++ [7] method was employed for CAM calculation, leveraging its ability to capture gradient information effectively and produce accurate heatmaps.
2.3 Visual Saliency Maps
Visual Saliency Maps [8] were generated using the SATSal model [16], a state-of-the-art technique for estimating visual saliency that showed promising results for both high and low-level scenes. Visual saliency refers to the perceptual prominence or importance of different regions within an image according to human perception and attention. By comparing CAMs with visual saliency maps, we aim to evaluate the alignment between model attention and human perception across different quantization levels.
2.4 Evaluation Metrics
To assess the similarity or difference between different CAMs at different quantization levels and between CAMs and visual saliency maps, we employed three metrics [17]: Similarity (SIM), Kullback-Leibler Divergence (KLD), and Pearson Correlation (CC).
Similarity (SIM) known as histogram intersection, is used to measure the similarity between two distributions. Introduced in [18] for the task of image matching.
KLD is a general information-theory-based metric. It measures the difference between two probability distributions.
CC, also known as the linear (Pearson) correlation coefficient. It is a metric used for measuring the dependency between two heatmaps. A high correlation between the maps means that the two distributions are closely related and interconnected.
2.5 Experimental Setup
We conducted experiments on a dataset consisting of $10,000$ images from the validation set of ImageNet[9]. Each neural network architecture was quantized at the specified levels using QAT, and CAMs were computed for each image in the dataset. Visual Saliency Maps were generated using the SATSal model for comparison with CAMs.
This comprehensive methodology enabled us to systematically investigate the impact of quantization on model perception and interpretability, providing insights into the behavior of neural networks across different quantization levels and their alignment with human perception.
Model Metric VGG-16 ResNet-50 DenseNet-121 MobileNet-V2 SqueezeNet-1.0 EfficientNet-B0
f32 v. GT SIM $\uparrow$ $0.485\pm 0.093$ $0.455\pm 0.089$ $0.473\pm 0.090$ $\textbf{0.506}\pm\textbf{0.087}$ $0.457\pm 0.090$ $0.487\pm 0.085$
CC $\uparrow$ $0.298\pm 0.227$ $0.260\pm 0.184$ $0.312\pm 0.193$ $\textbf{0.425}\pm\textbf{0.136}$ $0.259\pm 0.249$ $0.335\pm 0.219$
KLD $\downarrow$ $1.038\pm 0.419$ $1.091\pm 0.374$ $0.950\pm 0.363$ $\textbf{0.815}\pm\textbf{0.304}$ $1.057\pm 0.372$ $0.946\pm 0.337$
int16 v. GT SIM $\uparrow$ $0.703\pm 0.271$ $0.717\pm 0.276$ $0.733\pm 0.270$ $\textbf{0.741}\pm\textbf{0.255}$ $0.458\pm 0.091$ $0.717\pm 0.263$
CC $\uparrow$ $0.298\pm 0.229$ $0.255\pm 0.180$ $0.316\pm 0.194$ $\textbf{0.420}\pm\textbf{0.134}$ $0.258\pm 0.248$ $0.322\pm 0.218$
KLD $\downarrow$ $1.038\pm 0.421$ $1.065\pm 0.344$ $0.951\pm 0.373$ $\textbf{0.822}\pm\textbf{0.306}$ $1.054\pm 0.372$ $0.949\pm 0.337$
int8 v. GT SIM $\uparrow$ $0.694\pm 0.269$ $0.658\pm 0.273$ $0.636\pm 0.258$ $\textbf{0.704}\pm\textbf{0.253}$ $0.457\pm 0.091$ $0.679\pm 0.271$
CC $\uparrow$ $0.300\pm 0.231$ $0.247\pm 0.181$ $0.292\pm 0.197$ $\textbf{0.419}\pm\textbf{0.134}$ $0.257\pm 0.249$ $0.270\pm 0.218$
KLD $\downarrow$ $1.031\pm 0.413$ $1.095\pm 0.372$ $0.969\pm 0.371$ $\textbf{0.826}\pm\textbf{0.302}$ $1.054\pm 0.370$ $1.038\pm 0.348$
int16 v. f32 SIM $\uparrow$ $0.926\pm 0.085$ $\textbf{0.964}\pm\textbf{0.068}$ $0.963\pm 0.068$ $0.939\pm 0.062$ $0.926\pm 0.094$ $0.882\pm 0.114$
CC $\uparrow$ $0.878\pm 0.163$ $\textbf{0.939}\pm\textbf{0.168}$ $0.895\pm 0.223$ $0.832\pm 0.099$ $0.915\pm 0.184$ $0.751\pm 0.151$
KLD $\downarrow$ $0.075\pm 0.099$ $0.045\pm 0.125$ $\textbf{0.045}\pm\textbf{0.107}$ $0.050\pm 0.028$ $0.049\pm 0.105$ $0.176\pm 0.097$
int8 v. f32 SIM $\uparrow$ $\textbf{0.921}\pm\textbf{0.087}$ $0.844\pm 0.146$ $0.859\pm 0.111$ $0.916\pm 0.073$ $0.898\pm 0.089$ $0.807\pm 0.167$
CC $\uparrow$ $0.875\pm 0.161$ $0.609\pm 0.285$ $0.560\pm 0.264$ $0.755\pm 0.124$ $\textbf{0.893}\pm\textbf{0.196}$ $0.463\pm 0.235$
KLD $\downarrow$ $0.078\pm 0.098$ $0.289\pm 0.215$ $0.183\pm 0.123$ $0.075\pm 0.036$ $\textbf{0.064}\pm\textbf{0.111}$ $0.396\pm 0.172$
Table 1: Quantitative Results for models: GT represents the human visual saliency map, f32 represents the float 32 precision model, int16 represents the integer 16 precision model, int8 represents the integer 8 precision model
Fig. 2: Qualitative results of CAMs and visual saliency across multiple precisions and models.
3 Results
The obtained results shed light on the alterations in Class Activation Maps (CAMs) across various popular architectures when subjected to quantization. Our analysis revealed nuanced changes in model perception, indicating varying sensitivities of different architectures to quantization. The results of our experimentations are summarized in Table 1.
3.1 Empirical Results
The trend observed across the metrics indicates that int16 precision tends to perform relatively better than both f32 and int8 precision in terms of alignment with ground truth visual saliency maps.
The int16 precision models generally exhibit higher $SIM$ and $CC$ scores with ground truth visual saliency maps compared to both f32 and int8 precision models for most architectures. Additionally, Int 16 precision models often demonstrate lower $KLD$ values, suggesting better preservation of information between CAMs and ground truth visual saliency maps. We hypothesize that this discrepancy may be due to the preservation of noise in the model at higher precision levels, leading to a weaker alignment with human perception. On the other hand, the lower precision of int8 quantization may lead to more information loss or distortion in the CAMs, resulting in reduced alignment with human perception.
We also notice that $MobileNet$ consistently demonstrated robustness to quantization, maintaining high alignment between CAMs and Visual Saliency Maps across different precision levels. $SqueezeNet$ showed a consistent performance across all precision levels in comparison with visual saliency. These consistent results imply that it is less sensitive to quantization and maintains a stable alignment with ground truth visual saliency maps. This robustness could be advantageous in practical applications where maintaining interpretability and model perception is crucial, especially when deploying models on resource-constrained devices.
By establishing a reference point using human visual behavior through comparison with visual saliency maps, we gain valuable insights into how well the model’s attention aligns with human perception. However, it’s equally important to assess the relative changes in network perceptual behavior in comparison to the f32 baseline model.
Models like $ResNet50$ and $DenseNet121$ demonstrate relatively high SIM values, indicating substantial preservation of perceptual information during quantization to int16 precision. This observation suggests that these models maintain a similar visual interpretation capability across different precision levels. In contrast, $EfficientNet$ exhibits the most degradation in all metrics, indicating a comparatively lower degree of preservation of perceptual information under quantization.
The $VGG16$ model displays the highest preservation of similarity between int8 and f32 precision CAMs, indicating its effectiveness in retaining spatial details during quantization. Conversely, $SqueezeNet$ demonstrates the strongest linear correlation and the most aligned probability distributions between int8 and f32 precision CAMs, suggesting its robustness to quantization-induced changes in spatial perception.
These findings suggest that over-parameterized architectures like $VGG16$, $ResNet50$, and $DenseNet121$ exhibit strong robustness to quantization due to their high redundancy in parameters. In contrast, $SqueezeNet$ also demonstrates high robustness, which can be attributed to its efficient architecture and deliberate design choices aimed at maintaining performance under quantization.
3.2 Qualitative Results
Figure 2 depicts an example of an image from the study dataset with the related overlaid saliency map. It also depicts the different CAMs associated with different models and levels of precision.
$VGG-16$, $ResNet50$, and $DenseNet121$ exhibit a low level of impact in terms of changes on CAMs, at the int16 precision. On the other hand, we notice significant shifts when comparing at the int8 quantization level. Which is consistent with the deductions from the empirical results.
Smaller networks like $MobileNet$, and $SqueezeNet$ show fewer changes across lower levels. We can especially notice that CAMs from $SqueezeNet$ remain unchanged.
$EfficientNet$ displays the highest level of changes across the models which means that it is much less robust. This is due to the simplicity of the network architecture and the small number of parameters.
4 Conclusion
This study investigated how quantization affects different neural network architectures using Class Activation Maps (CAMs) to gauge their alignment with human perception. We found that quantization with int16 precision strikes a good balance between efficiency and maintaining human-like perception (measured by SIM, CC, KLD). Notably, MobileNet and SqueezeNet exhibited exceptional robustness due to their efficient designs. Conversely, EfficientNet suffered more under quantization.
These findings offer valuable insights for deploying neural networks on resource-constrained devices. By understanding how different architectures respond to quantization, researchers and practitioners can make informed decisions to achieve a balance between model performance, interpretability, and efficiency in real-world applications.

### v2

@@
 Abstract
-Neural network quantization is an essential technique for deploying models on resource-constrained devices.
-However, its impact on model perceptual fields, particularly regarding class activation maps (CAMs), remains a significant area of investigation.
-In this study, we explore how quantization alters the spatial recognition ability of the perceptual field of vision models, shedding light on the alignment between CAMs and visual saliency maps across various architectures.
-Leveraging a dataset of 10,000 images from ImageNet, we rigorously evaluate six diverse foundational CNNs: VGG16, ResNet50, EfficientNet, MobileNet, SqueezeNet, and DenseNet.
-We uncover nuanced changes in CAMs and their alignment with human visual saliency maps through systematic quantization techniques applied to these models.
-Our findings reveal the varying sensitivities of different architectures to quantization and underscore its implications for real-world applications in terms of model performance and interpretability.
-The primary contribution of this work revolves around deepening our understanding of neural network quantization, providing insights crucial for deploying efficient and interpretable models in practical settings.
-Index Terms—  Neural Network Quantization, Class Activation Maps (CAM), Model Interpretability, Visual Saliency
+Neural network quantization is a critical technique for deploying models on resource-limited devices.
+Despite its widespread use, the impact of quantization on model perceptual fields, particularly in relation to class activation maps (CAMs), remains underexplored.
+This study investigates how quantization influences the spatial recognition abilities of vision models by examining the alignment between CAMs and visual salient objects maps across various architectures.
+Utilizing a dataset of 10,000 images from ImageNet, we conduct a comprehensive evaluation of six diverse CNN architectures: VGG16, ResNet50, EfficientNet, MobileNet, SqueezeNet, and DenseNet.
+Through the systematic application of quantization techniques, we identify subtle changes in CAMs and their alignment with Salient object maps.
+Our results demonstrate the differing sensitivities of these architectures to quantization and highlight its implications for model performance and interpretability in real-world applications.
+This work primarily contributes to a deeper understanding of neural network quantization, offering insights essential for deploying efficient and interpretable models in practical settings.
+Index Terms—  Neural Network Quantization, Class Activation Maps (CAM), Model Interpretability, Salient objects detection
 1 Introduction
@@
 Quantization involves converting the weights and activations of a deep neural network from floating-point precision to lower bit-width representations.
-Quantization techniques all evolved from the following equation :
+Quantization techniques have evolved from the following equation:
 $Q(x)=\text{round}\left(\frac{x}{\Delta}\right)\times\Delta\vspace{-2mm}$ (1)
-where $x$ is the continuous value, $Q(x)$ presents the quantized value, and $\Delta$ is the quantization step size, determining the precision of the quantized representation.
+where $x$ is the continuous value, $Q(x)$ represents the quantized value, and $\Delta$ is the quantization step size, determining the precision of the quantized representation.
 By reducing the precision of parameters, quantization enables the delivery of lighter-weight models for efficient execution on devices with limited computational resources, such as smartphones [2], embedded systems [3], and IoT devices [4].
@@
 Ensuring that quantized models maintain interpretability is essential for building trust and facilitating human-machine collaboration.
-Class activation maps (CAMs) [5] are visualization techniques that highlight the regions of an image contributing most to a neural network’s decision.
+Class activation maps (CAMs) [5] are visualization techniques that highlight the regions of an image contributing most to a neural network’s output.
 CAMs provide valuable insights into where the model focuses its attention when making predictions, enhancing understanding and interoperability.
@@
 Visual saliency [8] refers to the perceptual prominence or importance of different regions within an image for the human visual system.
-Unlike CAMs, which are derived from the internal activations of a neural network, visual saliency is based on human perception and attention.
-Comparing CAMs with visual saliency maps [8] can help assess the alignment between model attention and human perception and evaluate the interpretability of quantized models.
-In this paper, we investigate the impact of quantization on model perception, focusing on CAMs and their alignment with visual saliency maps.
+Unlike CAMs, which are derived from the internal activations of a neural network, visual Saliency maps and salient objects maps is based on human perception and attention.
+Comparing CAMs with visual salient object maps [9] can help assess the alignment between model attention and human cognitive perception to evaluate the interpretability of quantized models.
+In this paper, we investigate the impact of quantization on model perception, focusing on CAMs and their alignment with visual salient objects maps.
 We evaluate six diverse neural network architectures on a comprehensive dataset of images from ImageNet [10].
 Through rigorous experimentation, we analyze the changes in CAMs induced by quantization and their implications for model interpretability and performance.
+Our findings contribute to advancing the understanding of neural network quantization and offer insights crucial for deploying efficient and interpretable models in practical settings.
 [^1] Source Code
-Our findings contribute to advancing the understanding of neural network quantization and offer insights crucial for deploying efficient and interpretable models in practical settings.
 2 Methodology
@@
 We selected six widely-used neural network architectures, namely VGG16 [11], ResNet50 [12], EfficientNet [13], MobileNet [14], SqueezeNet [15], and DenseNet [16], due to their diverse structures and widespread adoption across various computer vision tasks.
-Quantization, the focal subject of our study, was performed using the Quantization Aware Training (QAT) module from PyTorch 222 Quantization Pytorch only for inference, in order to simulate adaptive quantization.
+Quantization, the focal subject of our study, was performed using the Quantization Aware Training (QAT) module from PyTorch [^2] Quantization PyTorch only for inference, in order to simulate adaptive quantization.
 This technique allows for the quantization of neural network weights and activations while preserving the gradient flow necessary for calculating CAMs.
-Three Quantization levels were chosen: namely Float 32 (f32), Integer 16 (int16), and Integer 8 (int8).
-It’s worth noting that other quantization techniques in PyTorch do not support gradient flow during quantization.
+Three quantization levels were chosen: Float 32 (f32), Integer 16 (int16), and Integer 8 (int8).
+It is worth noting that other quantization techniques in PyTorch do not support gradient flow during quantization.
+In PyTorch, adaptive quantization is often implemented using Quantization-Aware Training (QAT), which involves simulating quantization effects during training while keeping the gradient flow intact as mentioned before.
+The adaptive quantization process in PyTorch can be described using the following equations:
+  1. 1.
+    Scale and Zero Point Calculation: For adaptive quantization, the scale $s$, which is a factor used to convert a floating-point value into an integer value, and zero point $z$, which is an integer value that maps the minimum floating-point value $x_{min}$ to the minimum quantized value $q_{min}$, are calculated based on the minimum and maximum values of the input tensor:
+    $s=\frac{x_{\text{max}}-x_{\text{min}}}{q_{\text{max}}-q_{\text{min}}}$ (2)
+    $z=q_{\text{min}}-\frac{x_{\text{min}}}{s}$ (3)
+    where $x_{\text{min}}$ and $x_{\text{max}}$ are the minimum and maximum values of the input tensor $x$.
+$q_{\text{min}}$ and $q_{\text{max}}$ are the minimum and maximum values of the quantized range (e.g., for int8, $q_{\text{min}}=0$ and $q_{\text{max}}=255$).
+  2. 2.
+    Quantization Equation: The quantization of a value $x$ is performed as follows:
+    $Q(x)=\text{round}\left(\frac{x}{s}+z\right)$ (4)
+Combining these steps, the full adaptive quantization process can be summarized:
+$Q(x)=\text{round}\left(\frac{x-x_{\text{min}}}{\frac{x_{\text{max}}-x_{\text{min}}}{q_{\text{max}}-q_{\text{min}}}}+q_{\text{min}}\right)$ (5)
 2.2 CAM Calculation
@@
 CAMs provide insights into the regions of an image that contribute most to a neural network’s decision-making process, facilitating interpretability and understanding of model behavior.
-The CamGrad++ [7] method was employed for CAM calculation, leveraging its ability to capture gradient information effectively and produce accurate heatmaps.
-2.3 Visual Saliency Maps
-Visual Saliency Maps [8] were generated using the SATSal model [16], a state-of-the-art technique for estimating visual saliency that showed promising results for both high and low-level scenes.
+The Grad-CAM++ [7] method was employed for CAM calculation, leveraging its ability to capture gradient information effectively and produce accurate heatmaps.
+2.3 Visual Salient Object Maps
+Visual Salient Object Maps were generated using the BiRefNet model [9], a state-of-the-art technique for detecting visual salient objects that showed promising results for both high and low-level scenes.
 Visual saliency refers to the perceptual prominence or importance of different regions within an image according to human perception and attention.
-By comparing CAMs with visual saliency maps, we aim to evaluate the alignment between model attention and human perception across different quantization levels.
+While visual salient objects are the set of objects that represent the most prominent items depicted in the scene, aiming to segment the foreground from the background of the scene.
+By comparing CAMs with visual salient objects maps, we aim to evaluate the alignment between model attention and human perception across different quantization levels.
 2.4 Evaluation Metrics
-To assess the similarity or difference between different CAMs at different quantization levels and between CAMs and visual saliency maps, we employed three metrics [17]: Similarity (SIM), Kullback-Leibler Divergence (KLD), and Pearson Correlation (CC).
-Similarity (SIM) known as histogram intersection, is used to measure the similarity between two distributions. Introduced in [18] for the task of image matching.
-KLD is a general information-theory-based metric. It measures the difference between two probability distributions.
-CC, also known as the linear (Pearson) correlation coefficient.
-It is a metric used for measuring the dependency between two heatmaps.
-A high correlation between the maps means that the two distributions are closely related and interconnected.
-2.5 Experimental Setup
-We conducted experiments on a dataset consisting of $10,000$ images from the validation set of ImageNet[9].
+To assess the similarity or difference between different CAMs at different quantization levels and between CAMs and visual salient objects maps, we employed three metrics [17]: Similarity (SIM), Kullback-Leibler Divergence (KLD), and Pearson Correlation (CC).
+Similarity (SIM): Known as histogram intersection, this metric is used to measure the similarity between two distributions. It was introduced in [18] for the task of image matching.
+[…]
+Where $S_{GT}$ represnets the salient object mask, and $S_{CAM}$ represents the Class Activation Map.
+Kullback-Leibler Divergence (KLD): This is a general information-theory-based metric that measures the difference between two probability distributions.
+$KL(S_{GT},S_{CAM})=\sum_{i}S_{CAMi}\log\left(\epsilon+\frac{S_{CAMi}}{\epsilon+S_{GTi}}\right)$ (7)
+Pearson Correlation (CC): Also known as the linear (Pearson) correlation coefficient, this metric is used for measuring the dependency between two heatmaps.
+A high correlation between the maps indicates that the two distributions are closely related and interconnected.
+$CC(S_{GT},S_{CAM})=\frac{Cov(S_{GT},S_{CAM})}{\sigma(S_{GT})\times\sigma(S_{CAM})}$ (8)
+We conducted experiments on a dataset consisting of 10,000 images randomly selected from the validation set of ImageNet [10].
 Each neural network architecture was quantized at the specified levels using QAT, and CAMs were computed for each image in the dataset.
-Visual Saliency Maps were generated using the SATSal model for comparison with CAMs.
 This comprehensive methodology enabled us to systematically investigate the impact of quantization on model perception and interpretability, providing insights into the behavior of neural networks across different quantization levels and their alignment with human perception.
 Model Metric VGG-16 ResNet-50 DenseNet-121 MobileNet-V2 SqueezeNet-1.0 EfficientNet-B0
-f32 v. GT SIM $\uparrow$ $0.485\pm 0.093$ $0.455\pm 0.089$ $0.473\pm 0.090$ $\textbf{0.506}\pm\textbf{0.087}$ $0.457\pm 0.090$ $0.487\pm 0.085$
-CC $\uparrow$ $0.298\pm 0.227$ $0.260\pm 0.184$ $0.312\pm 0.193$ $\textbf{0.425}\pm\textbf{0.136}$ $0.259\pm 0.249$ $0.335\pm 0.219$
-KLD $\downarrow$ $1.038\pm 0.419$ $1.091\pm 0.374$ $0.950\pm 0.363$ $\textbf{0.815}\pm\textbf{0.304}$ $1.057\pm 0.372$ $0.946\pm 0.337$
-int16 v. GT SIM $\uparrow$ $0.703\pm 0.271$ $0.717\pm 0.276$ $0.733\pm 0.270$ $\textbf{0.741}\pm\textbf{0.255}$ $0.458\pm 0.091$ $0.717\pm 0.263$
-CC $\uparrow$ $0.298\pm 0.229$ $0.255\pm 0.180$ $0.316\pm 0.194$ $\textbf{0.420}\pm\textbf{0.134}$ $0.258\pm 0.248$ $0.322\pm 0.218$
-KLD $\downarrow$ $1.038\pm 0.421$ $1.065\pm 0.344$ $0.951\pm 0.373$ $\textbf{0.822}\pm\textbf{0.306}$ $1.054\pm 0.372$ $0.949\pm 0.337$
-int8 v. GT SIM $\uparrow$ $0.694\pm 0.269$ $0.658\pm 0.273$ $0.636\pm 0.258$ $\textbf{0.704}\pm\textbf{0.253}$ $0.457\pm 0.091$ $0.679\pm 0.271$
-CC $\uparrow$ $0.300\pm 0.231$ $0.247\pm 0.181$ $0.292\pm 0.197$ $\textbf{0.419}\pm\textbf{0.134}$ $0.257\pm 0.249$ $0.270\pm 0.218$
-KLD $\downarrow$ $1.031\pm 0.413$ $1.095\pm 0.372$ $0.969\pm 0.371$ $\textbf{0.826}\pm\textbf{0.302}$ $1.054\pm 0.370$ $1.038\pm 0.348$
-int16 v. f32 SIM $\uparrow$ $0.926\pm 0.085$ $\textbf{0.964}\pm\textbf{0.068}$ $0.963\pm 0.068$ $0.939\pm 0.062$ $0.926\pm 0.094$ $0.882\pm 0.114$
-CC $\uparrow$ $0.878\pm 0.163$ $\textbf{0.939}\pm\textbf{0.168}$ $0.895\pm 0.223$ $0.832\pm 0.099$ $0.915\pm 0.184$ $0.751\pm 0.151$
-KLD $\downarrow$ $0.075\pm 0.099$ $0.045\pm 0.125$ $\textbf{0.045}\pm\textbf{0.107}$ $0.050\pm 0.028$ $0.049\pm 0.105$ $0.176\pm 0.097$
-int8 v. f32 SIM $\uparrow$ $\textbf{0.921}\pm\textbf{0.087}$ $0.844\pm 0.146$ $0.859\pm 0.111$ $0.916\pm 0.073$ $0.898\pm 0.089$ $0.807\pm 0.167$
-CC $\uparrow$ $0.875\pm 0.161$ $0.609\pm 0.285$ $0.560\pm 0.264$ $0.755\pm 0.124$ $\textbf{0.893}\pm\textbf{0.196}$ $0.463\pm 0.235$
-KLD $\downarrow$ $0.078\pm 0.098$ $0.289\pm 0.215$ $0.183\pm 0.123$ $0.075\pm 0.036$ $\textbf{0.064}\pm\textbf{0.111}$ $0.396\pm 0.172$
-Table 1: Quantitative Results for models: GT represents the human visual saliency map, f32 represents the float 32 precision model, int16 represents the integer 16 precision model, int8 represents the integer 8 precision model
-Fig. 2: Qualitative results of CAMs and visual saliency across multiple precisions and models.
+f32 v. GT SIM $\uparrow$ $0.44\pm 0.19$ $0.35\pm 0.17$ $0.37\pm 0.21$ $0.40\pm 0.22$ $0.43\pm 0.19$ $0.41\pm 0.20$
+CC $\uparrow$ $0.43\pm 0.27$ $0.22\pm 0.19$ $0.31\pm 0.22$ $0.40\pm 0.20$ $0.40\pm 0.28$ $0.37\pm 0.25$
+KLD $\downarrow$ $1.17\pm 0.89$ $1.53\pm 0.97$ $1.35\pm 0.96$ $1.24\pm 0.95$ $1.17\pm 0.89$ $1.29\pm 0.97$
+int16 v. GT SIM $\uparrow$ $0.67\pm 0.33$ $0.64\pm 0.36$ $0.67\pm 0.35$ $0.68\pm 0.34$ $0.43\pm 0.19$ $0.67\pm 0.33$
+CC $\uparrow$ $0.38\pm 0.27$ $0.21\pm 0.19$ $0.31\pm 0.22$ $0.40\pm 0.20$ $0.40\pm 0.29$ $0.36\pm 0.25$
+KLD $\downarrow$ $1.29\pm 0.97$ $1.63\pm 1.03$ $1.36\pm 0.94$ $1.22\pm 0.91$ $1.17\pm 0.89$ $1.32\pm 1.01$
+int8 v. GT SIM $\uparrow$ $0.66\pm 0.33$ $0.57\pm 0.36$ $0.56\pm 0.34$ $0.64\pm 0.34$ $0.43\pm 0.19$ $0.63\pm 0.34$
+CC $\uparrow$ $0.38\pm 0.28$ $0.20\pm 0.19$ $0.29\pm 0.22$ $0.40\pm 0.20$ $0.40\pm 0.29$ $0.30\pm 0.25$
+KLD $\downarrow$ $1.29\pm 0.97$ $1.65\pm 1.05$ $1.37\pm 0.95$ $1.24\pm 0.92$ $1.17\pm 0.90$ $1.40\pm 1.00$
+int16 v. f32 SIM $\uparrow$ $0.93\pm 0.09$ $0.96\pm 0.07$ $0.96\pm 0.07$ $0.94\pm 0.06$ $0.93\pm 0.09$ $0.88\pm 0.11$
+CC $\uparrow$ $0.88\pm 0.16$ $0.94\pm 0.17$ $0.90\pm 0.22$ $0.83\pm 0.10$ $0.92\pm 0.18$ $0.75\pm 0.15$
+KLD $\downarrow$ $0.08\pm 0.10$ $0.05\pm 0.13$ $0.05\pm 0.11$ $0.05\pm 0.03$ $0.05\pm 0.11$ $0.18\pm 0.10$
+int8 v. f32 SIM $\uparrow$ $0.92\pm 0.09$ $0.84\pm 0.15$ $0.86\pm 0.11$ $0.92\pm 0.07$ $0.90\pm 0.09$ $0.81\pm 0.17$
+CC $\uparrow$ $0.88\pm 0.16$ $0.61\pm 0.29$ $0.56\pm 0.26$ $0.76\pm 0.12$ $0.89\pm 0.20$ $0.46\pm 0.24$
+KLD $\downarrow$ $0.08\pm 0.10$ $0.29\pm 0.22$ $0.18\pm 0.12$ $0.08\pm 0.04$ $0.06\pm 0.11$ $0.40\pm 0.17$
+Table 1: Quantitative Results for models: GT represents the salient object map, f32 represents the float 32 precision model, int16 represents the integer 16 precision model, int8 represents the integer 8 precision model
+Fig. 2: Qualitative results of CAMs and visual Salient object across multiple precisions and models.
 3 Results
-The obtained results shed light on the alterations in Class Activation Maps (CAMs) across various popular architectures when subjected to quantization.
-Our analysis revealed nuanced changes in model perception, indicating varying sensitivities of different architectures to quantization.
-The results of our experimentations are summarized in Table 1.
+The results highlight significant alterations in Class Activation Maps (CAMs) across diverse popular architectures under quantization.
+Our analysis unveils subtle shifts in model perception, revealing varying sensitivities among different architectures to quantization effects.
+A summary of our findings is presented in Table 1.
 3.1 Empirical Results
-The trend observed across the metrics indicates that int16 precision tends to perform relatively better than both f32 and int8 precision in terms of alignment with ground truth visual saliency maps.
-The int16 precision models generally exhibit higher $SIM$ and $CC$ scores with ground truth visual saliency maps compared to both f32 and int8 precision models for most architectures.
-Additionally, Int 16 precision models often demonstrate lower $KLD$ values, suggesting better preservation of information between CAMs and ground truth visual saliency maps.
-We hypothesize that this discrepancy may be due to the preservation of noise in the model at higher precision levels, leading to a weaker alignment with human perception.
-On the other hand, the lower precision of int8 quantization may lead to more information loss or distortion in the CAMs, resulting in reduced alignment with human perception.
-We also notice that $MobileNet$ consistently demonstrated robustness to quantization, maintaining high alignment between CAMs and Visual Saliency Maps across different precision levels.
-$SqueezeNet$ showed a consistent performance across all precision levels in comparison with visual saliency.
-These consistent results imply that it is less sensitive to quantization and maintains a stable alignment with ground truth visual saliency maps.
-This robustness could be advantageous in practical applications where maintaining interpretability and model perception is crucial, especially when deploying models on resource-constrained devices.
-By establishing a reference point using human visual behavior through comparison with visual saliency maps, we gain valuable insights into how well the model’s attention aligns with human perception.
-However, it’s equally important to assess the relative changes in network perceptual behavior in comparison to the f32 baseline model.
-Models like $ResNet50$ and $DenseNet121$ demonstrate relatively high SIM values, indicating substantial preservation of perceptual information during quantization to int16 precision.
-This observation suggests that these models maintain a similar visual interpretation capability across different precision levels.
-In contrast, $EfficientNet$ exhibits the most degradation in all metrics, indicating a comparatively lower degree of preservation of perceptual information under quantization.
-The $VGG16$ model displays the highest preservation of similarity between int8 and f32 precision CAMs, indicating its effectiveness in retaining spatial details during quantization.
-Conversely, $SqueezeNet$ demonstrates the strongest linear correlation and the most aligned probability distributions between int8 and f32 precision CAMs, suggesting its robustness to quantization-induced changes in spatial perception.
-These findings suggest that over-parameterized architectures like $VGG16$, $ResNet50$, and $DenseNet121$ exhibit strong robustness to quantization due to their high redundancy in parameters.
-In contrast, $SqueezeNet$ also demonstrates high robustness, which can be attributed to its efficient architecture and deliberate design choices aimed at maintaining performance under quantization.
+The table 1 summarizes the performance metrics across various neural network architectures (VGG-16, ResNet-50, DenseNet-121, MobileNet-V2, SqueezeNet-1.0, EfficientNet-B0) under different quantization levels (f32, int16, int8).
+It evaluates the similarity (SIM), Pearson correlation coefficient (CC), and Kullback-Leibler divergence (KLD) metrics for each combination.
+Similarity (SIM): Across all architectures, quantization generally increases SIM compared to ground truth (GT), especially noticeable for int16 and int8 levels.
+VGG-16, ResNet-50, DenseNet-121, and MobileNet-V2 consistently show substantially higher SIM values compared to other models under the same quantization levels.
+The increase is due to the nature of the metric, as quantization may adjust the distributions of the salient object map and such that the minima operation in SIM captures more coinciding foreground regions between the two maps.
+SqueezeNet-1.0 shows high robustness in results.
+This can be attributed to the complex Fire modules that employs a squeeze and expansion operations to reduce and refine the amount of data.
+The model also bolsters complex skip connections that induce the propagation of information from shallow to deeper layers, decreasing loss of data, thus introducing an inherent robustness to information loss induced by methods for adaptive quantization.
+Pearson Correlation (CC): CC values vary significantly across architectures and quantization levels.
+Higher correlations (>0.5) are observed mainly between f32 and int16 levels, as evident in ResNet-50, DenseNet-121, and SqueezeNet-1.0.
+Lower correlation is observed in the quantization of EfficientNet-B0.
+This is attributed to the more simplistic architectural design relying on straightforward sequence of convolutional layers while minimizing the number of parameters ($5.3M$ parameters).
+This leads increased sensitivity compared to other model that employ more robust methods to propagate information from shallow to deep layers using skip connections (e.i.
+ResNet-50, DenseNet-121) or other models that use much larger number of parameters like VGG-16 ($138M$ parameters).
+The general trend indicates lower correlations at higher quantization levels, indicating greater divergence from ground truth maps.
+Kullback-Leibler Divergence (KLD): KLD values generally decrease with higher quantization precision (f32 to int8), suggesting improved alignment with ground truth Salient object maps.
+ResNet-50, DenseNet-121, and EfficientNet-B0 exhibit notable difference in KLD with increased quantization precision.
+These architectures are designed to be deeper making them more sensitive to information alteration induced by quantization during forward propagation especially at lower precision levels (i.e Int8) .
+Comparing int16 and int8 with f32 shows varying degrees of similarity and correlation changes across architectures.
+Generally, int16 maintains higher SIM and CC compared to int8, with exceptions observed in SqueezeNet-1.0.
+Results indicate that different neural network architectures react differently to quantization, with some models maintaining higher similarity and correlation with ground truth even under lower precision.
+Higher SIM and CC values suggest that certain architectures preserve visual interpretation capabilities better across different quantization levels, crucial for applications requiring accurate and reliable model interpretation.
+This indicates that Therefore, while the number of parameters can be indicative of model complexity and potentially its sensitivity to quantization, the actual sensitivity also depends on the interplay of architectural design, feature representation, quantization methodology, and how these factors interact with each other.
 3.2 Qualitative Results
-Figure 2 depicts an example of an image from the study dataset with the related overlaid saliency map. It also depicts the different CAMs associated with different models and levels of precision.
-$VGG-16$, $ResNet50$, and $DenseNet121$ exhibit a low level of impact in terms of changes on CAMs, at the int16 precision.
-On the other hand, we notice significant shifts when comparing at the int8 quantization level.
-Which is consistent with the deductions from the empirical results.
-Smaller networks like $MobileNet$, and $SqueezeNet$ show fewer changes across lower levels. We can especially notice that CAMs from $SqueezeNet$ remain unchanged.
-$EfficientNet$ displays the highest level of changes across the models which means that it is much less robust.
-This is due to the simplicity of the network architecture and the small number of parameters.
+Figure 2 depicts the impact of quantization on CAMs for the chosen deep learning models.
+Each row represents a different model (DenseNet-121, EfficientNet-B0, MobileNetV2, ResNet-50, SqueezeNet-1.0, VGG-16), and each column represents the effect of a different quantization level (f32, int16, int8).
+The first column shows the original image and the second a ground truth segmentation of the salient object.
+VGG-16 demonstrates a gradual shift in the perceptual field, where with the decrease of the quantization level, it expands its perceptual field, thus introducing more information noise about other parts of the image.
+ResNEt-50 shows more robustness at the int 16 quantization level, but a significant change in the perceptual field areas at the int8 level.
+DenseNet-121 shows similar behavior as ResNEt-50, but we notice that it tends to use a larger perceptual field covering more area of the stimuli.
+MobileNet-v2 shows less robustness to quantization effects, where the perceptual field changes at all precision levels.
+We also notice that the network tends to have a larger and more uniform perceptual field.
+It also tends to heavily emphasise on larger areas of the image as shown by the heatmap.
+In accordance with the empirical data, SqueezeNet-1.0 is the least affected by the quantization, we can notice that it tends to have uniforme and more moderate coverage of the image maximizing the amount of innformation gathered, this explains it’s hardiness against the loss of information stemming from quantization.
+The EfficientNet-B0 network shows the least robustness in accordance with the empirical data. The shift in the perceptual fields across different quantization levels is evident through the figure.
+In summary, our comprehensive analysis reveals the profound impact of quantization on the perceptual fields of various neural network architectures, where both quantitative and qualitative analysis align their findings.
 4 Conclusion
-This study investigated how quantization affects different neural network architectures using Class Activation Maps (CAMs) to gauge their alignment with human perception.
-We found that quantization with int16 precision strikes a good balance between efficiency and maintaining human-like perception (measured by SIM, CC, KLD).
-Notably, MobileNet and SqueezeNet exhibited exceptional robustness due to their efficient designs.
-Conversely, EfficientNet suffered more under quantization.
+This study investigated how quantization affects different neural network architectures using Class Activation Maps (CAMs) to gauge their alignment with human perception and the semantic content of the stimulus.
+The study demonstrates that while some models, such as SqueezeNet-1.0, exhibit remarkable robustness to quantization, others, like EfficientNet-B0, show significant sensitivity, particularly at lower precision levels.
+These findings underscore the importance of considering quantization effects when deploying deep learning models on resource-constrained devices especially in use cases of interpretability.
 These findings offer valuable insights for deploying neural networks on resource-constrained devices.
 By understanding how different architectures respond to quantization, researchers and practitioners can make informed decisions to achieve a balance between model performance, interpretability, and efficiency in real-world applications.
+Nonetheless, we need to mention that these results are still limited and that it would be beneficial to expand these results to more diverse datasets, models, domains, and tasks.